OpenStudy (anonymous):

PLEASE HELP: A barrel contains 160 gallons of water and is being drained at a constant rate of 7 gallons per hour. Write an equation that models the number of gallons, g, after t hours.

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Initial contents = 160 gallons. This is a simple rate problem: v = rt, where v is the volume drained, r is the rate of draining (7 gal/hr) , and t is the amount of time that has passed. Your equation will take the initial contents and subtract the amount that has drained. What do you think it is?

OpenStudy (anonymous):

22.85? @whpalmer4

OpenStudy (whpalmer4):

Is 22.85 an equation modeling the number of gallons, g, after t hours?

OpenStudy (anonymous):

huh?????? @whpalmer4

OpenStudy (whpalmer4):

You have to read the question carefully, and answer what it asks. It doesn't ask how many hours it will take to drain the barrel.

OpenStudy (whpalmer4):

You did a great job of answering that question (well, you didn't include the unit), but that wasn't what they asked for!

OpenStudy (anonymous):

do i just write an equation? @whpalmer4

OpenStudy (whpalmer4):

What part of "Write an equation that models the number of gallons, g, after t hours." isn't making sense to you?

OpenStudy (anonymous):

is it 160=7t @whpalmer4

OpenStudy (whpalmer4):

No. Look, the amount in the barrel is 160 gallons at t = 0, right? After that, some number of gallons drains out, the amount being found by v = rt where r = 7 gal/hr and t is the number of hours.

OpenStudy (whpalmer4):

160 = 7t makes no sense when you plug in a number of hours. 160 = 7 * 1 hours? Fails both numerically and dimensionally.

OpenStudy (anonymous):

i still dont understand=(...... im sorry im just really tired it 12 in the morning so im not understanding this as much @whpalmer4

OpenStudy (whpalmer4):

A valid equation will have valid units for the results. For example, the v = rt equation has gallons on the left = gallons/hour * hours on the right, so we have gallons = gallons

OpenStudy (whpalmer4):

Amount in the barrel at the start is 160 gallons. So at t =0, our equation to show how much remains in the barrel is simply: W = 160 gallons. If we let water drain out for 1 hour, the barrel will lose (7 gallons/hr)(1 hr) = 7 gallons and will now contain 160-7 gallons = 153 gallons. After another hour, the barrel will have lost (7 gallons/hr)(2 hr) = 14 gallons and will now contain 160-14 = 146 gallons. Do you see the pattern? After we open the drain, and 7 gallons/hour starts pouring out, the barrel will contain W = 160 gallons - (7 gallons/hour)t where t is in hours. Let's check it out: at t = 0, W = 160 gallons - (7 gallons/hour)(0 hours) = 160 gallons at t = 22.85 hours W = 160 gallons - (7 gallons/hour)(22.85 hours) = 0 gallons

OpenStudy (anonymous):

so the equation is W= 160-(7g/h)(0h)=160? @whpalmer4

OpenStudy (whpalmer4):

No. Do you see a variable called t in there anywhere? That line you copied is an evaluation of the function at the time t = 0. The equation is W = 160 gallons - (7 gallons/hr)*t My suggestion is that you get some sleep, and look at it again in the morning.

OpenStudy (anonymous):

ok thank you and sorry for not being able to understand the problem, Thank you for your help=) @whpalmer4

OpenStudy (whpalmer4):

You're welcome. Tomorrow's another day...fresh eyes often bring fresh insight.

hero (hero):

g(t) = 160 - 7t

hero (hero):

t = number of hours of constant drainage

OpenStudy (anonymous):

Thank you very much hero @Hero

hero (hero):

No need to torture the students @whpalmer4